OpenStudy (anonymous):

Yale would like to use the completing the square method, while Kiara is considering using the factoring method to find exact solutions of 3x2 - 6x = 17. Using complete sentences, explain who has chose the better method and why.

OpenStudy (anonymous):

Kiara needs to use the following steps: 3x^2 - 6x - 17 = 0 Which does not factor nicely. So Yale's method would be better: x^2 - 2x = 17/3 x^2 - 2x + 1 = 17/3 + 1 = 20/3 (x - 1)^2 = 20/3 x - 1 = +- sqrt(20/3) x = 1 +- sqrt(20/3) = 1 +- 2sqrt(15)/3 = (3 +- 2sqrt 15)/3